OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of √180 feet and a width of √125 feet.

OpenStudy (mathmagician):

\[P=10\sqrt{5}+12\sqrt{5}=22\sqrt{5}\]

Directrix (directrix):

= 72 √5 + 10√5 = 82√ 5 ---> perimeter

OpenStudy (mathmagician):

\[\sqrt{180}=\sqrt{6*6*5}=6\sqrt{5}\]

OpenStudy (mathmagician):

not \[36\sqrt{5}\]

Directrix (directrix):

Thanks.